OpenStudy (anonymous):

Solve x – 2y = 6 3x – 6y = 18 A. (2, –2) B. (3, negative three halves) C. No Solutions D. Infinitely Many Solutions

OpenStudy (anonymous):

solve for x using the first equation x-2y=6 x=6+2y Substitute in for x in second equation 3(6+2y)-6y=18 18+6y-6y=18 18=18 Infinitely Many Solutions

OpenStudy (anonymous):

Thankyou so much!!!!! :)

OpenStudy (nincompoop):

i hope you understand what has to be done next time